An object is projected with a velocity U at an angle R to the horizontal.It just clears two walls of equal height b and at a distance of b apart. Show that 2U cosR√(U² sin²R - 2bg) = bg?

1
Expert's answer
2022-03-07T18:06:33-0500

Explanations & Calculations


  • If the vertical component of the velocity at the first time the trajectory marks its height b is "\small \uparrow V" then by applying "\small v^2=u^2+2as" for the vertical motion from start to that point we can link those moments.

"\qquad\qquad
\begin{aligned}
\small V^2&=\small (U\sin R)^2+2(-g)b\\
\small V&=\small \sqrt{U^2\sin^2 R-2gb}
\end{aligned}"

  • Then we can calculate the time it takes to travel between the two pillars of height b or the distance in between them; b. Make note the objects horizontal velocity is a constant throughout which is "\small U\cos R."

"\qquad\qquad
\begin{aligned}
\small \to\\
\small s&=\small ut\\
\small b&=\small U\cos R.t\\
\small t&=\small \frac{b}{U\cos R}
\end{aligned}"

  • Then, during this time it travels vertically as well but makes a 0 vertical displacement as a whole.
  • By this happening vertically,

"\qquad\qquad
\begin{aligned}
\small \uparrow\\
\small s&=\small ut+1/2at^2\\
\small 0&=\small Vt+\frac{1}{2}(-g)t^2\\
\small t&=\small \frac{2V}{g}
\end{aligned}"


  • Finally, equalling those two obtains via time and through a little bit of arrangement, we get what we need.

"\qquad\qquad
\begin{aligned}
\small \frac{2V}{g}&=\small \frac{b}{U\cos R}\\
\small 2U\cos R.V&=\small gb\\
\small bg&=\small 2U\cos R.\sqrt{U^2\sin^2 R-2gb}
\end{aligned}"